# Dexterous Manipulation Policies from RGB Human Videos via 4D Hand-Object Trajectory Reconstruction

### v1

@@ IV Experiments
We conduct two sets of experiments with our approach: grasping of objects and manipulating different objects. We present the main results in Fig. 2. Video rollouts for the results are best viewed in our website.
@@ IV-A Grasping Experiments
Setup. We collect 20 human videos of daily object grasping, with one video per object; the object categories are shown in Fig. 4 (a). Each video lasts 3–6 seconds and is recorded at 30 fps. We train our models across all objects and evaluate grasp success in the IsaacGym simulator using an 18-Dof Inspire robotic hand, as its size closely matches that of a human hand. The object meshes and poses in the simulator are initialized using the reconstructed meshes and poses obtained from the recorded videos. Each grasp $q^{\text{grasp}}\in\mathbb{R}^{18}$ is subjected to a 300-step disturbance phase in IsaacGym, during which forces are applied sequentially from six directions: $\pm x$, $\pm y$, and $\pm z$. The applied force magnitude is set to 0.5 times the object’s mass. A grasp is considered successful if the object’s displacement under applied force disturbances remains within 3 cm of its initial position. Through these experiments, we aim to answer the following questions:
@@ IV-B Manipulation Experiments
A3: Our VideoManip captures hand-object motion from human videos, providing spatiotemporal supervision for manipulation policy learning. The reconstructed 4D trajectories closely align with the corresponding RGB human videos, as shown in Fig. 5. Leveraging both these reconstructed trajectories and additional synthesized trajectories generated by DemoGen, the trained DP3 policy achieves the highest overall success rate of 62.86% across all tasks (Tab. II). The synthesized trajectories enable our policy to generalize towards different target object locations(videos available on our website). While, DemoGen based trajectory augmentation is less helpful for tasks like Close Drawer and Screw Bulb, where objects (Bulb and Socket) are tightly coupled, the variations in object locations are still helpful for generalization of the trained policy.
@@
Figure 5: Visualization of VideoManip Execution Using In-Scene (top) and In-the-Wild (bottom) Video Data Sources. For each task, given RGB human videos (row 1), we reconstruct 4D trajectories of the human hand and objects (row 2). Trained with these trajectories, we executed on a real-world LEAP Hand (row 3).
A4: The Pour Tea task is evaluated on both in-scene and in-the-wild video categories using $^{\text{world}}T_{\text{cam}}$ and $^{\text{gra}}R_{\text{cam}}$, respectively. Table II shows no significant performance difference between the two video sources; however, without applying $^{\text{gra}}R_{\text{cam}}$ for the in-the-wild Pour Tea task, the success rate drops to 0%, highlighting the importance of gravity-alignment calibration. We further ablate the effect of DemoGen on the in-scene Pour Tea task by adding $x$-$y$ plane perturbations within [0.2,0.2][-0.2,0.2],m to the target object Bowl position reconstructed from the source video, and synthesizing between 1 and 1000 trajectories. During evaluation, the Bowl is placed at 15 different locations within this perturbed region. As the number of synthesized trajectories increases, the success rate improves steadily from 1/15 to 13/15 (Fig. 2(d)), demonstrating that DemoGen enables DP3 to learn actions that generalize across target object locations.
@@ V Conclusion, Limitations, and Future Work
We present a reconstruction-based framework that recovers 4D hand–object trajectories from RGB-only, in-the-wild human videos at low cost, without relying on specialized sensing devices. The reconstructed trajectories are shown to be effective for training dexterous grasping and manipulation policies, and provide more physically grounded supervision compared to retargeting-based approaches that rely on generated videos.
Despite these advantages, our method has several limitations. (1) Our framework relies on multiple off-the-shelf 3D vision models, including image-to-mesh reconstruction and object pose estimation, which may introduce errors that can compound across stages. We currently circumvent this issue by recording the human videos from approximately egocentric viewpoints that facilitate reliable hand and object reconstruction, instead of downloading generic videos from the internet. In future work, we plan to incorporate a trajectory verification module to automatically detect and filter erroneous reconstructions, and scale to more in-the-wild videos. (2) The current framework assumes static or approximately static camera setups. We plan to extend our approach to dynamic camera settings by leveraging recent advances in dynamic scene reconstruction [60, 61]. (3) Our manipulation policy learning currently relies on synthesized trajectories from DemoGen represented as 3D point clouds. In real-world manipulation, however, object point clouds are often difficult to track due to occlusions caused by the robot hand. We currently assume that the hand–object relative contact established during grasping is preserved throughout execution to update the object point clouds along with robot hand point clouds. In future work, we plan to explore alternative trajectory synthesis modalities beyond point clouds, such as image-based representations.

### v2

@@ IV Experiments
We conduct two sets of experiments with our framework: grasping of objects and manipulating different objects. We present the main results in Fig. 2. Video rollouts for the results are best viewed in our website.
@@ IV-A Grasping Experiments
Setup. We collect 20 videos of humans grasping objects, with one video per object; the object categories are shown in Fig. 4 (a). Each 30 fps video lasts 3 to 6 seconds. We train our models across all objects and evaluate grasp success in the IsaacGym simulator using an 18-Dof Inspire robot hand, as its size closely matches that of a human hand. We initialize the simulator with object meshes and poses reconstructed from the recorded videos. Each grasp $q^{\text{grasp}}\in\mathbb{R}^{18}$ undergoes a 300-step disturbance phase in IsaacGym, with forces applied sequentially from six directions: $\pm x$, $\pm y$, and $\pm z$. The applied force magnitude is set to 0.5 times the object’s mass. We consider a grasp successful if the object’s displacement under applied force disturbances stays within 3 cm of its initial position. Through these experiments, we aim to answer the following questions:
@@ IV-B Manipulation Experiments
A3: VideoManip captures hand-object motion from human videos, providing spatiotemporal supervision for manipulation policy learning. The reconstructed trajectories closely align with the corresponding RGB human videos, as shown in Fig. 5. Leveraging both these reconstructed trajectories and additional synthesized trajectories generated by DemoGen, the trained DP3 policy achieves the highest overall success rate of 62.86% across all tasks (Tab. II). The synthesized trajectories enable our policy to generalize towards different target object locations(videos available on our website). While, DemoGen based trajectory augmentation is less helpful for tasks like Close Drawer and Screw Bulb, where objects (Bulb and Socket) are tightly coupled, the variations in object locations are still helpful for generalization of the trained policy.
@@
Figure 5: Visualization of VideoManip Execution Using In-Scene (top) and In-the-Wild (bottom) Video Data Sources. For each task, given RGB human videos (row 1), we reconstruct 3D trajectories of the human hand and objects (row 2). Trained with these trajectories, we executed on a real-world LEAP Hand (row 3).
A4: The Pour Tea task is evaluated on both in-scene and in-the-wild video categories using $^{\text{world}}T_{\text{cam}}$ and $^{\text{gra}}R_{\text{cam}}$, respectively. Table II shows no significant performance difference between the two video sources; however, without applying $^{\text{gra}}R_{\text{cam}}$ for the in-the-wild Pour Tea task, the success rate drops to 0%, highlighting the importance of gravity-alignment calibration. We further ablate the effect of DemoGen on the in-scene Pour Tea task by adding $x$-$y$ plane perturbations within [0.2,0.2][-0.2,0.2] m to the target object Bowl position reconstructed from the source video, and synthesizing between 1 and 1000 trajectories. During evaluation, the Bowl is placed at 15 different locations within this perturbed region. As the number of synthesized trajectories increases, the success rate improves steadily from 1/15 to 13/15 (Fig. 2(d)), demonstrating that DemoGen enables DP3 to learn actions that generalize across target object locations.
@@ V Conclusion, Limitations, and Future Work
We present a reconstruction-based framework that recovers 3D hand–object trajectories from RGB-only, in-the-wild human videos at low cost, without relying on specialized sensing devices. The reconstructed trajectories are shown to be effective for training dexterous grasping and manipulation policies, and provide more physically grounded supervision compared to retargeting-based approaches that rely on generated videos.
Despite these advantages, VideoManip has several limitations. (1) VideoManip framework relies on multiple 3D vision models, including image-to-mesh reconstruction and object pose estimation, which may introduce compounding errors across stages. We currently circumvent this issue by recording the human videos from approximately egocentric viewpoints that facilitate reliable hand and object reconstruction, instead of downloading generic videos from the internet. In future work, we plan to incorporate a trajectory verification module to automatically detect and filter erroneous reconstructions, and scale to more in-the-wild videos. (2) The current framework assumes static or approximately static camera setups. We plan to extend our approach to dynamic camera settings by leveraging recent advances in dynamic scene reconstruction [60, 61]. (3) The manipulation policy learning currently relies on synthesized trajectories from DemoGen represented as 3D point clouds. In real-world manipulation, however, object point clouds are often difficult to track due to occlusions caused by the robot hand. We currently assume that the hand–object relative contact established during grasping is preserved throughout execution to update the object point clouds along with robot hand point clouds. In future work, we plan to explore alternative trajectory synthesis modalities beyond point clouds, such as image-based representations.